# Spec-Gaussian: Anisotropic View-Dependent Appearance for 3D Gaussian Splatting

### v1

@@ Abstract
The recent advancements in 3D Gaussian splatting (3D-GS) have not only facilitated real-time rendering through modern GPU rasterization pipelines but have also attained state-of-the-art rendering quality. Nevertheless, despite its exceptional rendering quality and performance on standard datasets, 3D-GS frequently encounters difficulties in accurately modeling specular and anisotropic components. This issue stems from the limited ability of spherical harmonics (SH) to represent high-frequency information. To overcome this challenge, we introduce Spec-Gaussian, an approach that utilizes an anisotropic spherical Gaussian (ASG) appearance field instead of SH for modeling the view-dependent appearance of each 3D Gaussian. Additionally, we have developed a coarse-to-fine training strategy to improve learning efficiency and eliminate floaters caused by overfitting in real-world scenes. Our experimental results demonstrate that our method surpasses existing approaches in terms of rendering quality. Thanks to ASG, we have significantly improved the ability of 3D-GS to model scenes with specular and anisotropic components without increasing the number of 3D Gaussians. This improvement extends the applicability of 3D GS to handle intricate scenarios with specular and anisotropic surfaces. Our codes and datasets will be released.
Figure 1: Our method not only achieves real-time rendering but also significantly enhances the capability of 3D-GS to model scenes with specular and anisotropic components. Key to this enhanced performance is our use of ASG appearance field to model the appearance of each 3D Gaussian, which results in substantial improvements in rendering quality for both complex and general scenes. Moreover, we employ anchor Gaussians to constrain the geometry of point-based representations, thereby improving the ability of 3D-GS to accurately model reflective parts and accelerating both training and rendering processes.
@@ 1 Introduction
High-quality reconstruction and photorealistic rendering from a collection of images are crucial for a variety of applications, such as augmented reality/virtual reality (AR/VR), 3D content production, and art creation. Classic methods employ primitive representations, like meshes [34] and points [4, 60], and take advantage of the rasterization pipeline optimized for contemporary GPUs to achieve real-time rendering. In contrast, neural radiance fields (NeRF) [32, 6, 33] utilize neural implicit representation to offer a continuous scene representation and employ volumetric rendering to produce rendering results. This approach allows for enhanced preservation of scene details and more effective reconstruction of scene geometries.
@@
Despite its exceptional performance, 3D-GS struggles to model specular components within scenes (see Fig. 1). This issue primarily stems from the limited ability of low-order spherical harmonics (SH) to capture the high-frequency information required in these scenarios. Consequently, this poses a challenge for 3D-GS to model scenes with reflections and specular components, as illustrated in Fig. 1 and Fig. 7.
To address the issue, we introduce a novel approach called Spec-Gaussian, which combines anisotropic spherical Gaussian (ASG) [54] for modeling anisotropic and specular components, anchor-based geometry-aware 3D Gaussians for acceleration and storage reduction, and an effective training mechanism to eliminate floaters and improve learning efficiencies. Specifically, the method incorporates three key designs: 1) A new 3D Gaussian representation that utilizes an ASG appearance field instead of SH to model the appearance of each 3D Gaussian. ASG with a few orders can effectively model high-frequency information that low-order SH cannot. This new design enables 3D-GS to more effectively model anisotropic and specular components in static scenes. 2) A hybrid approach employing sparse anchor points to control the location and representation of its child Gaussians. This strategy results in a hierarchical and geometry-aware point-based scene representation and enables us to store only the anchor Gaussians, significantly reducing storage requirements and enhancing the geometry. 3) A coarse-to-fine training scheme specifically tailored for 3D-GS is designed to eliminate floaters and boost learning efficiency. This strategy effectively shortens learning time by optimizing low-resolution rendering in the initial stage, preventing the need to increase the number of 3D Gaussians and regularizing the learning process to avoid the generation of unnecessary geometric structures that lead to floaters.
@@
  • A novel ASG appearance field to model the view-dependent appearance of each 3D Gaussian, which enables 3D-GS to effectively represent scenes with specular and anisotropic components without sacrificing rendering speed.
  • An anchor-based hybrid model to reduce the computational and storage overhead brought by learning the ASG appearance field.
@@
  • An anisotropic dataset has been made to assess the capability of our model in representing anisotropy. Extensive experiments show the effectiveness of our method in modeling scenes with specular highlights and anisotropy.
@@ 2.1 Implicit Neural Radiance Fields
Neural rendering has attracted significant interest in the academic community for its unparalleled ability to generate photorealistic images. Methods like NeRF [32] utilize Multi-Layer Perceptrons (MLPs) to model the geometry and radiance fields of a scene. Leveraging the volumetric rendering equation and the inherent continuity and smoothness of MLPs, NeRF achieves high-quality scene reconstruction from a set of posed images, establishing itself as the state-of-the-art (SOTA) method for novel view synthesis. Subsequent research has extended the utility of NeRF to various applications, including mesh reconstruction [46, 25, 52], inverse rendering [42, 63, 29, 56], optimization of camera parameters [27, 48, 47, 36], few-shot learning [12, 55, 51], and anti-aliasing [2, 1, 3].
@@
To reduce the training time of MLP-based NeRF methods and improve rendering speed, subsequent work has enhanced NeRF’s efficiency in various ways. Structure-based techniques [61, 13, 38, 16, 8] have sought to improve inference or training efficiency by caching or distilling the implicit neural representation into more efficient data structures. Hybrid methods [28, 43] increase efficiency by incorporating explicit voxel-based data structures. Factorization methods [5, 17, 9, 15] apply a low-rank tensor assumption to decompose the scene into low-dimensional planes or vectors, achieving better geometric consistency. Compared to continuous implicit representations, the convergence of individual voxels in the grid is independent, significantly reducing training time. Additionally, Instant-NGP [33] utilizes a hash grid with a corresponding CUDA implementation for faster feature querying, enabling rapid training and interactive rendering of neural radiance fields.
@@
Figure 2: Pipeline of our proposed Spec-Gaussian. The optimization process begins with SfM points derived from COLMAP or generated randomly, serving as the initial state for the anchor Gaussians. Within a view frustum, $k$ neural Gaussians are spawned from each visible anchor Gaussian, using the corresponding offsets. Their other attributes, such as opacity $\sigma$, rotation $r$, and scaling $s$, are decoded through the respective tiny MLPs. To address the limitations of low-order SH and pure MLP in modeling high-frequency information, we additionally employ ASG in conjunction with a feature decoupling MLP to model the view-dependent appearance of each neural Gaussian. Subsequently, neural Gaussians with opacity $\sigma>0$ are rendered through a differentiable Gaussian rasterization pipeline, effectively capturing specular highlights and anisotropy in the scene.
@@ 2.2 Point-based Neural Radiance Fields
Recently, 3D-GS [21] has employed anisotropic 3D Gaussians, initialized from Structure from Motion (SfM), to represent 3D scenes. The innovative densification mechanism and CUDA-customized differentiable Gaussian rasterization pipeline of 3D-GS have not only achieved state-of-the-art (SOTA) rendering quality but also significantly surpassed the threshold of real-time rendering. Many concurrent works have rapidly extended 3D-GS to a variety of downstream applications, including dynamic scenes [31, 57, 58, 18, 24, 50], text-to-3D generation [26, 44, 10, 59, 11], avatars [65, 64, 19, 39, 35], and scene editing [53, 7].
Despite achieving SOTA results on commonly used benchmark datasets, 3D-GS still struggles to model scenes with specular and reflective components, which limits its practical application in real-time rendering at the photorealistic level. In this work, by replacing spherical harmonics (SH) with an anisotropic spherical Gaussian (ASG) appearance field, we have enabled 3D-GS to model complex specular scenes more effectively. Furthermore, this improvement enhances rendering quality in general scenes without significantly impacting rendering speed.
@@ 3 Method
The overview of our method is illustrated in Fig. 2. The input to our model is a set of posed images of a static scene, together with a sparse point cloud obtained from SfM [40]. The core of our method is to use the ASG appearance field to replace SH in modeling the appearance of 3D Gaussians (Sec. 3.2). To reduce the storage overhead and rendering speed pressure introduced by ASG, we design a hybrid Gaussian model that employs sparse anchor Gaussians to facilitate the generation of neural Gaussians (Sec. 3.3) to model the 3D scene. Finally, we introduce a simple yet effective coarse-to-fine training strategy to reduce floaters in real-world scenes (Sec. 3.4).
@@ 3.1 Preliminaries
3.1.1 3D Gaussian Splatting
3D-GS [21] is a point-based method that employs anisotropic 3D Gaussians to represent scenes. Each 3D Gaussian is defined by a center position $\bm{x}$, opacity $\sigma$, and a 3D covariance matrix $\Sigma$, which is decomposed into a quaternion $\bm{r}$ and scaling $\bm{s}$. The view-dependent appearance of each 3D Gaussian is represented using the first three orders of spherical harmonics (SH). This method not only retains the rendering details offered by volumetric rendering but also achieves real-time rendering through a CUDA-customized differentiable Gaussian rasterization process. Following [66], the 3D Gaussians can be projected to 2D using the 2D covariance matrix $\Sigma^{\prime}$, defined as:
@@
where $J$ is the Jacobian of the affine approximation of the projective transformation, and $V$ represents the view matrix, transitioning from world to camera coordinates. To facilitate learning, the 3D covariance matrix $\Sigma$ is decomposed into two learnable components: the quaternion $\bm{r}$, representing rotation, and the 3D-vector $\bm{s}$, representing scaling. The resulting $\Sigma$ is thus represented as the combination of a rotation matrix $R$ and scaling matrix $S$ as:
@@
$C(\textbf{p})=\sum_{i\in N}T_{i}\alpha_{i}c_{i},\quad\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\textbf{p}-\mu_{i})^{T}\sum^{\prime}(\textbf{p}-\mu_{i})},$ (3)
@@
3.1.2 Anisotropic Spherical Gaussian.
Anisotropic spherical Gaussian (ASG) [54] has been designed within the traditional rendering pipeline to efficiently approximate lighting and shading. Different from spherical Gaussian (SG), ASG has been demonstrated to effectively represent anisotropic scenes with a relatively small number. In addition to retaining the fundamental properties of SG, ASG also exhibits rotational invariance and can represent full-frequency signals. The ASG function is defined as:
@@ 3.2 Anisotropic View-Dependent Appearance
3.2.1 ASG Appearance Field for 3D Gaussians.
@@
Although ASG theoretically enhance the ability of SH to model anisotropy, directly using ASG to represent the specular color of each 3D Gaussian still falls short in accurately modeling anisotropic and specular components, as demonstrated in Fig. 5. Inspired by [15], we do not use ASG directly to represent color but instead employ ASG to model the latent feature of each 3D Gaussian. This latent feature, containing anisotropic information, is then fed into a tiny feature decoupling MLP $\Psi$ to determine the final specular color:
@@
where $\kappa$ is the latent feature derived from ASG, $\bigoplus$ denotes the concatenation operation, $\gamma$ represents the positional encoding, $\mathbf{d}$ is the unit view direction pointing from the camera to each 3D Gaussian, $n$ is the normal of each 3D Gaussian that will be discussed in Sec. 3.2.2, and $\omega_{r}$ is the unit reflect direction. This strategy significantly enhances the ability of 3D-GS to model scenes with complex optical phenomena, whereas neither pure ASG nor pure MLP can achieve anisotropic appearance modeling as effectively as our approach.
3.2.2 Normal Estimation
Directly estimating the normals of 3D Gaussians presents a challenge, as 3D-GS comprises a collection of discrete entities, each representing a local space within a certain range, without forming a continuous surface. The calculation of normals typically necessitates a continuous surface, and the anisotropic shape of each entity in 3D-GS further complicates the determination of normals. Following [20, 41], we use the shortest axis of each Gaussian as its normal. This approach is based on the observation that 3D Gaussians tend to flatten gradually during the optimization process, allowing the shortest axis to serve as a reasonable approximation for the normal.
@@ 3.2.2 Normal Estimation
Figure 3: Using a coarse-to-fine strategy, our approach is able to optimize the scene in a progressive manner and eliminate the floaters efficiently.
3.3 Anchor-Based Gaussian Splatting
3.3.1 Neural Gaussian Derivation with ASG Appearance Field.
While the ASG appearance field significantly improves the ability of 3D-GS to model specular and anisotropic features, it introduces additional storage and computational overhead compared to using pure SH due to the additional local features $\mathbf{f}$ associated with each Gaussian. Although real-time rendering at over 100 FPS is still achievable in bounded scenes, the substantial increase in storage overhead and reduction in rendering speed caused by ASG in real-world unbounded scenes is unacceptable. Inspired by [30], we employ anchor-based Gaussian splatting to reduce storage overhead and the number of 3D Gaussians required for rendering, thereby accelerating the rendering.
Tri-MipRF 3D-GS Scaffold-GS GS-Shader Ours GT
Figure 4: Visualization on NeRF dataset. Our method has successfully achieved local specular highlights modeling, a capability that other 3D-GS-based methods fail to accomplish, while maintaining fast rendering speed. Compared to Tri-MipRF, a NeRF-based method, we have significantly enhanced the ability to model anisotropic materials.
Unlike the attributes carried by each entity in 3D-GS, each anchor Gaussian carries a position coordinate $\mathbf{P}_{v}\in\mathbb{R}^{3}$, a local feature $\mathbf{f}_{v}\in\mathbb{R}^{32}$, a displacement factor $\eta_{v}\in\mathbb{R}^{3}$, and $k$ learnable offsets $\mathbf{O}_{v}\in\mathbb{R}^{k\times 3}$. We use the sparse point cloud obtained from COLMAP [40] to initialize each anchor 3D Gaussian, serving as the voxel centers to guide the generation of neural Gaussians. The position $\mathbf{P}_{v}$ of the anchor Gaussian is initialized as:
@@ 3.3 Anchor-Based Gaussian Splatting
We then use the anchor Gaussians to guide the generation of neural Gaussians, which have the same attributes as vanilla 3D-GS. For each visible anchor Gaussian within the viewing frustum, we spawn $k$ neural Gaussians and predict their attributes (see Fig. 2). The positions $\mathbf{x}$ of neural Gaussians are calculated as:

### v2

@@ Abstract
The recent advancements in 3D Gaussian splatting (3D-GS) have not only facilitated real-time rendering through modern GPU rasterization pipelines but have also attained state-of-the-art rendering quality. Nevertheless, despite its exceptional rendering quality and performance on standard datasets, 3D-GS frequently encounters difficulties in accurately modeling specular and anisotropic components. This issue stems from the limited ability of spherical harmonics (SH) to represent high-frequency information. To overcome this challenge, we introduce Spec-Gaussian, an approach that utilizes an anisotropic spherical Gaussian (ASG) appearance field instead of SH for modeling the view-dependent appearance of each 3D Gaussian. Additionally, we have developed a coarse-to-fine training strategy to improve learning efficiency and eliminate floaters caused by overfitting in real-world scenes. Our experimental results demonstrate that our method surpasses existing approaches in terms of rendering quality. Thanks to ASG, we have significantly improved the ability of 3D-GS to model scenes with specular and anisotropic components without increasing the number of 3D Gaussians. This improvement extends the applicability of 3D GS to handle intricate scenarios with specular and anisotropic surfaces. Our codes and datasets are available at https://ingra14m.github.io/Spec-Gaussian-website.
@@ 1 Introduction
High-quality reconstruction and photorealistic rendering from a collection of images are crucial for a variety of applications, such as augmented reality/virtual reality (AR/VR), 3D content production, and art creation. Classic methods employ primitive representations, like meshes [36] and points [4, 62], and take advantage of the rasterization pipeline optimized for contemporary GPUs to achieve real-time rendering. In contrast, neural radiance fields (NeRF) [34, 5, 35] utilize neural implicit representation to offer a continuous scene representation and employ volumetric rendering to produce rendering results. This approach allows for enhanced preservation of scene details and more effective reconstruction of scene geometries. $\dagger$ Corresponding Authors.
@@
Despite its exceptional performance, 3D-GS struggles to model specular components within scenes (see Fig. 1). This issue primarily stems from the limited ability of low-order spherical harmonics (SH) to capture the high-frequency information in these scenarios. Consequently, this poses a challenge for 3D-GS to model scenes with reflections and specular components, as illustrated in Fig. 1.
Figure 1: Our method not only achieves real-time rendering but also significantly enhances the capability of 3D-GS to model scenes with specular and anisotropic components. Key to this enhanced performance is our use of ASG appearance field to model the appearance of each 3D Gaussian, which results in substantial improvements in rendering quality for both complex and general scenes.
To address the issue, we introduce a novel approach called Spec-Gaussian, which combines anisotropic spherical Gaussian (ASG) [55] for modeling anisotropic and specular components, an effective training mechanism to eliminate floaters and improve learning efficiencies, and anchor-based 3D Gaussians for acceleration and storage reduction. Specifically, the method incorporates two key designs: 1) A new 3D Gaussian representation that utilizes an ASG appearance field instead of SH to model the appearance of each 3D Gaussian. ASG with a few orders can effectively model high-frequency information that low-order SH cannot. This new design enables 3D-GS to more effectively model anisotropic and specular components in static scenes. 2) A coarse-to-fine training scheme specifically tailored for 3D-GS is designed to eliminate floaters and boost learning efficiency. This strategy effectively shortens learning time by optimizing low-resolution rendering in the initial stage, preventing the need to increase the number of 3D Gaussians and regularizing the learning process to avoid the generation of unnecessary geometric structures that lead to floaters.
@@
  • A novel ASG appearance field to model the view-dependent appearance of each 3D Gaussian, which enables 3D-GS to effectively represent scenes with specular and anisotropic components.
@@
  • An anisotropic dataset has also been made to assess the capability of our model in representing anisotropy.
@@ 2.1 Implicit Neural Radiance Fields
Neural rendering has attracted significant interest in the academic community for its unparalleled ability to generate photorealistic images. Methods like NeRF [34] utilize Multi-Layer Perceptrons (MLPs) to model the geometry and radiance fields of a scene. Leveraging the volumetric rendering equation and the inherent continuity and smoothness of MLPs, NeRF achieves high-quality scene reconstruction from a set of posed images, establishing itself as the state-of-the-art (SOTA) method for novel view synthesis. Subsequent research has extended the utility of NeRF to various applications, including mesh reconstruction [48, 24, 53, 31], inverse rendering [44, 66, 28, 57], optimization of camera parameters [26, 50, 49, 38], few-shot learning [11, 56, 52], and anti-aliasing [2, 1, 3].
@@
To reduce the training time of MLP-based NeRF methods and improve rendering speed, subsequent work has enhanced NeRF’s efficiency in various ways. Structure-based techniques [63, 12, 40, 15, 7] have sought to improve inference or training efficiency by caching or distilling the implicit neural representation into more efficient data structures. Hybrid methods [27, 45] increase efficiency by incorporating explicit voxel-based data structures. Factorization methods [5, 16, 8, 14] apply a low-rank tensor assumption to decompose the scene into low-dimensional planes or vectors, achieving better geometric consistency. Compared to continuous implicit representations, the convergence of individual voxels in the grid is independent, significantly reducing training time. Additionally, Instant-NGP [35] utilizes a hash grid with a corresponding CUDA implementation for faster feature querying, enabling rapid training and interactive rendering of neural radiance fields. Spec-NeRF [32] achieves high-quality specular reflection modeling by introducing Gaussian directional encoding.
@@
Figure 2: Pipeline of Spec-Gaussian. The optimization process begins with SfM points derived from COLMAP or generated randomly, serving as the initial state for the 3D Gaussians. To address the limitations of low-order SH and pure MLP in modeling high-frequency information, we additionally employ ASG in conjunction with a feature decoupling MLP to model the view-dependent appearance of each 3D Gaussian. Then, 3D Gaussians with opacity $\sigma>0$ are rendered through a differentiable Gaussian rasterization pipeline, effectively capturing specular highlights and anisotropy in the scene.
@@ 2.2 Point-based Neural Radiance Fields
Recently, 3D-GS [20] has employed anisotropic 3D Gaussians, initialized from Structure from Motion (SfM), to represent 3D scenes. The innovative densification mechanism and CUDA-customized differentiable Gaussian rasterization pipeline of 3D-GS have not only achieved state-of-the-art (SOTA) rendering quality but also significantly surpassed the threshold of real-time rendering. Many concurrent works have rapidly extended 3D-GS to a variety of downstream applications, including dynamic scenes [30, 58, 59, 17, 23], text-to-3D generation [25, 46, 9, 61, 10], avatars [68, 67, 18, 41, 37], scene editing [54, 6], and quality enhancement [33].
Despite achieving SOTA results on commonly used benchmark datasets, 3D-GS still struggles to model scenes with specular and reflective components, which limits its practical application in real-time rendering at the photorealistic level. In this work, by replacing spherical harmonics (SH) with an anisotropic spherical Gaussian (ASG) appearance field, we have enabled 3D-GS to model complex specular scenes more effectively.
@@ 3 Method
The overview of our method is illustrated in Fig. 2. The input to our model is a set of posed images of a static scene, together with a sparse point cloud obtained from SfM [42]. The core of our method is to use the ASG appearance field to replace SH in modeling the appearance of 3D Gaussians (Sec. 3.2). Moreover, we introduce a simple yet effective coarse-to-fine training strategy to reduce floaters in real-world scenes (Sec. 3.3). To further reduce the storage overhead and rendering speed pressure introduced by ASG, we combine a hybrid Gaussian model that employs sparse anchor Gaussians to facilitate the generation of neural Gaussians (Sec. 3.4) to model the 3D scene.
@@ 3.1 Preliminaries
3D Gaussian splatting.
3D-GS [20] is a point-based method that employs anisotropic 3D Gaussians to represent scenes. Each 3D Gaussian is defined by a center position $\boldsymbol{x}$, opacity $\sigma$, and a 3D covariance matrix $\Sigma$, which is decomposed into a quaternion $\boldsymbol{r}$ and scaling $\boldsymbol{s}$. The view-dependent appearance of each 3D Gaussian is represented using the first three orders of spherical harmonics (SH). This method not only retains the rendering details offered by volumetric rendering but also achieves real-time rendering through a CUDA-customized differentiable Gaussian rasterization process. Following [69], the 3D Gaussians can be projected to 2D using the 2D covariance matrix $\Sigma^{\prime}$, defined as:
@@
where $J$ is the Jacobian of the affine approximation of the projective transformation, and $V$ represents the view matrix, transitioning from world to camera coordinates. To facilitate learning, the 3D covariance matrix $\Sigma$ is decomposed into two learnable components: the quaternion $\boldsymbol{r}$, representing rotation, and the 3D-vector $\boldsymbol{s}$, representing scaling. The resulting $\Sigma$ is thus represented as the combination of a rotation matrix $R$ and scaling matrix $S$ as:
@@
$C(\textbf{p})=\sum_{i\in N}T_{i}\alpha_{i}c_{i},\quad\alpha_{i}=\sigma_{i}e^{-\frac{1}{2}(\textbf{p}-\mu_{i})^{T}\sum^{-1}(\textbf{p}-\mu_{i})},$ (3)
@@
Anisotropic spherical Gaussian.
Anisotropic spherical Gaussian (ASG) [55] has been designed in the traditional rendering pipeline to efficiently approximate lighting and shading. Different from spherical Gaussian (SG), ASG has been demonstrated to effectively represent anisotropic scenes with a small number. In addition to retaining the fundamental properties of SG, ASG also exhibits rotational invariance and can represent full-frequency signals. The ASG function is defined as:
@@
Anchor-based Gaussian splatting.
Anchor-based Gaussian splatting was first proposed by Scaffold-GS [29]. Unlike the attributes carried by each entity in 3D-GS, each anchor Gaussian carries a position coordinate $\mathbf{P}_{v}\in\mathbb{R}^{3}$, a local feature $\mathbf{f}_{v}\in\mathbb{R}^{32}$, a displacement factor $\eta_{v}\in\mathbb{R}^{3}$, and $k$ learnable offsets $\mathbf{O}_{v}\in\mathbb{R}^{k\times 3}$. They use the COLMAP [42] point cloud to initialize each anchor 3D Gaussian, serving as the voxel centers to guide the generation of neural Gaussians. The position $\mathbf{P}_{v}$ of the anchor Gaussian is initialized as:
[…]
Then anchor Gaussians can guide the generation of neural Gaussians, which have the same attributes as vanilla 3D-GS. For each visible anchor Gaussian within the viewing frustum, we spawn $k$ neural Gaussians and predict their attributes. The positions $\mathbf{x}$ of neural Gaussians are calculated as:
@@ 3.2 Anisotropic View-Dependent Appearance
ASG appearance field for 3D Gaussians.
@@
Although ASG theoretically enhance the ability of SH to model anisotropy, directly using ASG to represent the specular color of each 3D Gaussian still falls short in accurately modeling anisotropic and specular components, as demonstrated in Fig. 9. Inspired by [14], we do not use ASG directly to represent color but instead employ ASG to model the latent feature of each 3D Gaussian. This latent feature, containing anisotropic information, is then fed into a tiny feature decoding MLP $\Psi$ to determine the final specular color:
@@
where $\kappa$ is the latent feature derived from ASG, $\bigoplus$ denotes the concatenation operation, $\gamma$ represents the positional encoding, $\mathbf{d}$ is the unit view direction pointing from the camera to each 3D Gaussian, $n$ is the normal of each 3D Gaussian, and $\omega_{r}$ is the unit reflect direction. This strategy significantly enhances the ability of 3D-GS to model scenes with complex optical phenomena, whereas neither pure ASG nor pure MLP can achieve anisotropic appearance modeling as effectively as our approach.
Normal estimation.
Following [19, 43], we use the shortest axis of each Gaussian as its normal. This approach is based on the observation that 3D Gaussians tend to flatten gradually during the optimization process, allowing the shortest axis to serve as a reasonable approximation for the normal.
@@
[…]
[7.5cm]
Dataset Anisotropic Synthetic
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ FPS Mem Num.(k)
3D-GS 33.82 0.966 0.062 325 47MB 201
Scaffold-GS 35.34 0.972 0.052 234 27MB -
Ours-w/ anchor 36.76 0.976 0.046 180 25MB -
Ours-light 37.42 0.979 0.044 159 45MB 146
Ours 37.70 0.980 0.042 145 57MB 183
Figure 3: Quantitative Comparison on anisotropic synthetic dataset.
[…]
[5.8cm]Refer to caption
Figure 4: Using a coarse-to-fine strategy, our approach can eliminate the floaters without increasing the number of GS.
3.3 Coarse-to-fine Training
We observed that in many real-world scenarios, 3D-GS tends to overfit the training data, leading to the emergence of numerous floaters when rendering images from novel viewpoints. One important reason is that the COLMAP point cloud is too sparse. Poor initialization makes it difficult for 3D-GS to compensate for overly sparse areas through densification during optimization, leading to floaters in the rendering images. Moreover, 3D-GS accumulates gradients from each pixel to the GS: $\frac{dL}{d\mathbf{x}}=\sum\frac{dL}{d\mathbf{p}_{i}}\frac{d\mathbf{p}_{i}}{d\mathbf{x}}$, and the densification occurs when the accumulated amount exceeds a threshold $\tau_{g}=0.0002$. However, having positive and negative gradients can cause GSs that should be densified to be ignored due to the large negative gradient.
Thus, to mitigate the occurrence of floaters in real-world scenes, we propose a coarse-to-fine training mechanism. We first impose an L1 constraint on the gradients from pixels to GS: $\frac{dL}{d\mathbf{x}}=\sum\|\frac{dL}{d\mathbf{p}_{i}}\frac{d\mathbf{p}_{i}}{d\mathbf{x}}\|_{1}$, accumulating the numerical contribution from pixels to GS rather than gradients. This idea is similar to the concurrent works [60, 64]. Next, to avoid overfitting caused by excessive growth of 3D-GS during the early stages of optimization, we decide to train 3D-GS progressively from low to high resolution in real-world scenes:
$r(i)=\min(\lfloor r_{s}+(r_{e}-r_{s})\cdot i/{\tau}\rceil,r_{e}),$ (12)
where $r(i)$ is the image resolution at the $i$-th training iteration, $r_{s}$ is the starting image resolution, $r_{e}$ is the ending image resolution (the full resolution we aim to render), and $\tau$ is the threshold iteration, empirically set to 5k.
This training method allows 3D-GS to densify correctly and prevents excessive growth of 3D-GS in the early stages. Additionally, due to the lower resolution training in the initial phase, this mechanism reduces training time by approximately 10%. In our experiments, we offer a performance version with $\tau_{g}=0.0005$ and light version with $\tau_{g}=0.0006$.
3.4 Adaption for Anchor-Based Gaussian Splatting
While the ASG appearance field significantly improves the ability of 3D-GS to model specular and anisotropic features, it introduces additional computational overhead due to the additional local features $\mathbf{f}$ associated with each Gaussian. Inspired by [29], we employ anchor-based Gaussian splatting to reduce storage overhead and accelerate the rendering.
Figure 5: Visualization on NeRF dataset. Our method has achieved specular highlights modeling, which other 3D-GS-based methods fail to accomplish, while maintaining fast rendering speed.